### v1

@@ Abstract
In this study, we investigate the $P$-$V$ criticality and Ruppeiner geometry in the extended phase space of Hayward anti-de Sitter (AdS) black holes. Through thermodynamic analysis, we confirm that Hayward-AdS black holes undergo distinct $P$-$V$ phase transitions and exhibit well-defined critical phenomena in the vicinity of their critical points. These behaviors are characterized by four critical exponents that typically obey the scaling laws predicted by mean-field theory–indicating a consistent thermodynamic framework with classical phase transition systems ($e.g.,$ van der Waals fluids). Furthermore, we employ Ruppeiner geometry to probe the thermodynamic fluctuations of Hayward-AdS black holes, and by calculating the corresponding curvature scalar, we gain direct insights into the interaction nature of the black hole’s microscopic constituents.
@@ I Introduction
Pioneering studies by Hawking and Bekenstein (Bekenstein1972, ; Bardeen1973, ; Hawking:1976de, ) revealed the fundamental nature of black holes—not merely gravitational systems, but entities with distinct thermodynamic properties. Specifically, a black hole’s temperature is determined by its surface gravity, while its entropy is directly associated with the area of its event horizon (Bekenstein1973, ). Crucially, the Hawking temperature of a black hole changes as it accretes matter or emits radiation (Hawking:1974sw, ); this thermodynamic behavior strongly implies that black holes must possess a microscopic structure Strominger:1996sh . Since the prediction of this property, one of the most compelling questions in black hole physics has persisted: what exactly constitutes the microscopic state of a black hole?
Over the past few decades, the field of black hole thermodynamics has witnessed numerous groundbreaking discoveries. Similar to ordinary thermodynamic systems, black holes exhibit a rich variety of phase transitions—among which, those of black holes in anti-de Sitter (AdS) spacetimes have emerged as a focal point of recent research. A core assumption underpinning these studies is the interpretation of the cosmological constant $\Lambda$ as a thermodynamic pressure variable (Kastor:2009wy, ; Dolan:2010ha, ; Kubiznak:2016qmn, ), defined as
@@
where $L$ denotes the radius of the $D$-dimensional AdS spacetime and $G_{\rm N}$ is Newton’s gravitational constant. Within the framework of this definition, a variety of distinct phase transitions emerge in black hole thermodynamics, such as van der Waals Kubiznak:2012wp , (multiple) re-entrant Altamirano:2013ane ; Frassino:2014pha ; Dehyadegari:2017hvd ; Hendi:2017mfu , those with isolated critical points Dolan:2014vba , those with triple points Altamirano:2013uqa , and superfluid-like phase transitions Hennigar:2016xwd . (See Refs.Bhattacharya:2017hfj ; Spallucci:2013osa ; Majhi:2016txt ; Altamirano:2014tva ; Cai:2013qga ; Abdusattar:2023xxs ; Hu:2018qsy for more details on related works and extended analyses of these phase transitions.) [^1] Notably, recent studies have revealed a captivating finding: in black hole systems with quantum conformal anomaly, the scaling laws can be violated under certain conditions Hu:2024ldp ; Abdusattar:2024zzi .
@@
By analyzing the phase transitions of black holes, thermodynamic geometry has emerged as a powerful tool for probing their microscopic structure, with Ruppeiner geometry playing a particularly pivotal role (Ruppeinerb2008, ). The core idea of this approach is to define a line element in the thermodynamic parameter space, which quantifies the “distance” between two adjacent fluctuation states (Ruppeiner95, ). The curvature scalar constructed from this line element directly indicates the nature of interactions among the system’s components: a positive curvature corresponds to repulsive interactions, while a negative curvature signifies attractive interactions. Notably, this interpretation is not unique to black holes, but is derived from the application of Ruppeiner geometry to ordinary thermodynamic systems (Janyszek1990, ; Oshima1999, ), whose cross-system consistency lends reliability to microscopic studies of black holes. Inspired by this, the microscopic origin of black hole thermodynamics has been explored Wei:2012ui ; Wei:2015iwa ; Wei:2019uqg ; Wei:2019yvs ; Guo:2019oad ; Miao:2017fqg ; KordZangeneh:2017lgs ; Wei:2019ctz ; Dehyadegari:2016nkd ; Hu:2020pmr ; Abdusattar:2023xxs . It is important to note that the statistical mechanical study of black holes differs fundamentally from that of ordinary thermodynamic systems: while ordinary systems typically derive macroscopic thermodynamic quantities from the behavior of microscopic particles, black hole research requires “reverse deduction”—inferring information about microscopic structures from known thermodynamic quantities by analyzing macroscopic phase transition behavior.
@@
This paper focuses on the Hayward-AdS black hole, investigating its microstructure through a $P$-$V$ criticality analysis. A well-recognized limitation of classical gravity theories is the existence of spacetime singularities; quantum effects are widely believed to remedy this defect, highlighting the necessity of developing a quantum theory of gravity. However, no mature quantum gravity theory currently exists, making regular black hole models within a semiclassical framework an important avenue for studying singularity correction. Bardeen was the first to propose a singularity-free regular black hole model (bardeen1968non, ), which replaces the singularity of traditional black holes with a de Sitter core. Subsequent studies further demonstrated that the charged version of the Bardeen regular black hole can be derived from Einstein gravity coupled to a nonlinear electrodynamic source (AyonBeato:1998ub, ; AyonBeato:2000zs, ), and can even be interpreted as the gravitational field generated by a nonlinear magnetic monopole. Hayward proposed another representative class of regular black hole solutions (Hayward:2005gi, ): similar to the Bardeen solution, it is a degenerate configuration of the gravitational field of a nonlinear source, carrying a magnetic charge and a free integration constant. In theories with infinite towers of higher-order curvature corrections in dimensions $D\geq 5$, regular black holes with anti-de Sitter asymptotics are constructed Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025ftm ; these black holes, along with related branes and solutions coupled to various electrodynamics, are analyzed regarding regularity, core properties, and thermodynamics, where thermodynamic features linked to regularization parameters are also revealed. Motivated by this work, this paper focuses on analyzing the phase structure of the Hayward-AdS black hole and exploring the repulsive interactions among its microscopic components.
@@ II Thermodynamics of the Hayward-AdS black holes
The action for the Hayward-AdS black hole in dimensions $D\geq 5$ is given by Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025ftm
@@
where $g$ denotes the metric determinant, and $R$ is the Ricci scalar curvature, ${\alpha}_{n}$ are arbitrary coupling constants with dimensions of length2(n-1). The densities $\mathscr{Z}_{n}$ are selected by the condition that they possess second-order equations on general spherically symmetric ansätze. Varying the action in Eq.(2) yields the field equations, from which the line element of the static, spherically symmetric Hayward-AdS black hole is given by Hennigar:2025ftm
@@
Here ${\rm d}\Omega^{2}_{D-2}$ represents the line element of a (D2)(D-2)-dimensional sphere, $m$ is an integration constant that is related to the Arnowitt–Deser–Misner (ADM) mass or thermodynamic mass $M$ Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025ftm
@@
where $\Omega_{D-2}$ is the volume of a unit sphere $2\pi^{(D-1)/2}/\Gamma\left[(D-1)/2\right]$. The event horizon of this black hole is determined by the condition $f(r_{+})=0$, using which the explicit expression for mass $M$ can be written as Hennigar:2025ftm
$M=\frac{(D-2)\Omega_{D-2}r_{+}^{D-1}}{16\pi G_{\rm N}L^{2}}\left(1+\frac{L^{2}}{r_{+}^{2}-\alpha}\right)\,.$ (6)
The Hawking temperature of the Hayward-AdS black hole on an event horizon is obtained as Hennigar:2025ftm
@@
the first law of black hole thermodynamics and the corresponding Smarr formula take the form, respectively Dolan:2014vba ; Hennigar:2025ftm
@@ III.1 $P$-$V$ phase transition and critical behavior
By rearranging the expression in Eq. (7) and combining it with Eq. (1), one can derive the thermodynamic equation of state for the $D$-dimensional spherically symmetric Hayward-AdS black hole, which is given by Hennigar:2025ftm
@@
has a critical point solution $T=T_{\rm c},\ P=P_{\rm c},\ V=V_{\rm c}$. This system is known to exhibit a small-large black hole phase transition Hennigar:2025ftm , where the critical point
@@
The corresponding $\widetilde{P}-$$\widetilde{V}$ phase diagram is presented in Fig. 1.
Figure 1: $\widetilde{P}-$$\widetilde{V}$ phase diagram of Hayward-AdS black hole. For $\widetilde{T}>1$, the phase diagram exhibits ideal gas behavior, indicating that the black hole exists in a single unique phase. For $\widetilde{T}<1$, it shows oscillatory behavior, revealing the presence of two distinct phases (small black hole and large black hole). These two phases undergo a first-order phase transition, which terminates at the critical point $(\widetilde{V},\widetilde{P})=(1,1)$. Here we set $\alpha=0.2,G_{\rm N}=1.$
@@ III.2 Critical exponents near the critical point
where $t=\widetilde{T}-1$, $\omega=\widetilde{V}-1$. These two reduced parameters characterize the critical behavior of the system in the vicinity of the critical point. Here, $s$ and $l$ correspond to the horizon radii of the small and large black holes, respectively, which are the endpoints of the black hole’s coexistence phase. It is straightforward to verify that these critical exponents satisfy the following thermodynamic scaling laws Kubiznak:2012wp
Since the black hole entropy is also determined by its horizon radius, $C_{V}$ (constant-volume heat capacity) vanishes-implying the first critical exponent $\widetilde{\alpha}=0$. To calculate the remaining three critical exponents, it is convenient to expand the equation of state given in Eq. (12) around the critical point
@@
$a_{10}$ $=$ $\frac{(D-2){T_{c}}\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{3}{D-1}}}{4G_{\rm N}{P_{c}}\left\{\alpha-\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{2}{D-1}}\right\}^{2}}\,$, (22)
$a_{11}$ $=$ $-\frac{\left\{(D-2){T_{c}}\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{3}{D-1}}\left[3\alpha+\left(\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right)^{\frac{2}{D-1}}\right]\right\}}{4\left\{(D-1)G_{\rm N}{P_{c}}\left[\left(\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right)^{\frac{2}{D-1}}-\alpha\right]^{3}\right\}}\,$,\,\,
@@
$-D\left(D(D(D(D(41-19D)+114)-333)+189)+27\sqrt{3}\sqrt{(D-2)D}\right)$
$\Theta$ $=$ $\left(2\sqrt{3}\sqrt{(D-2)D^{5}}+3D^{3}-27D-9\sqrt{3}\sqrt{(D-2)D}+27\right)$
all of which are not zero. From Eq.(21), we derive $d\widetilde{P}=(a_{11}t+3a_{03}\omega^{2})d\omega$, with $t$ constant along the isothermal curve. Below the critical temperature ($i.e.,$ $T<T_{c}$), the black hole system exhibits two coexisting phases: the small black hole phase and the large black hole phase. We thus denote the solutions corresponding to these phases as ($\omega_{s}$, $\widetilde{V}_{s}$) (small black hole) and ($\omega_{l}$, $\widetilde{V}_{l}$) (large black hole).
Substituting this into the Maxwell equal-area law Spallucci:2013osa ; Majhi:2016txt ; Bhattacharya:2017hfj ; Lan:2015bia -given by ${P}^{*}\left(\widetilde{V}_{s}-\widetilde{V}_{l}\right)=\int^{s}_{l}\widetilde{P}d\widetilde{V}$, we further obtain
@@
where ${P}^{*}$ denotes the reduced pressure corresponding to the straight line connecting the two coexisting phases. Additionally, the endpoints of the small and large black hole phases (analogous to the vapor and liquid phases in classical systems) share the same pressure $i.e.,$ ${P}^{*}=\widetilde{P}_{l}=\widetilde{P}_{s}$ Spallucci:2013osa ; Majhi:2016txt ; Bhattacharya:2017hfj ; Lan:2015bia . This implies
@@
To conclude, we have derived four critical exponents for the Hayward-AdS black hole, which are given by
@@
In particular, these exponents are identical to those of the van der Waals liquid-gas system and thus satisfy the thermodynamic scaling laws (see Eq.III.2).

### v2

@@ Abstract
We investigate $P$-$V$ criticality, with a focus on Ruppeiner geometry, in the extended phase space of Hayward anti-de Sitter (AdS) black holes. Through thermodynamic analysis, we confirm that Hayward-AdS black holes undergo distinct $P$-$V$ phase transitions and exhibit well-defined critical phenomena in the vicinity of their critical points. These behaviors are characterized by four critical exponents that typically obey the scaling laws predicted by mean-field theory–indicating a consistent thermodynamic framework with classical phase transition systems ($e.g.,$ van der Waals fluids). Furthermore, we employ Ruppeiner geometry to probe the thermodynamic fluctuations of Hayward-AdS black holes, and by calculating the corresponding curvature scalar, we gain direct insights into the interaction nature of the black hole’s microscopic constituents.
@@ I Introduction
Pioneering studies by Hawking and Bekenstein (Bekenstein1972, ; Bardeen1973, ; Hawking:1976de, ) revealed the fundamental nature of black holes–not merely gravitational systems, but entities with distinct thermodynamic properties. Specifically, a black hole’s temperature is determined by its surface gravity, while its entropy is directly associated with the area of its event horizon (Bekenstein1973, ). Crucially, the Hawking temperature of a black hole changes as it accretes matter or emits radiation (Hawking:1974sw, ); this thermodynamic behavior strongly implies that black holes must possess a microscopic structure Strominger:1996sh . Since the prediction of this property, one of the most compelling questions in black hole physics has persisted: what exactly constitutes the microscopic state of a black hole?
Over the past few decades, the field of black hole thermodynamics has witnessed numerous groundbreaking discoveries. Similar to ordinary thermodynamic systems, black holes exhibit a rich variety of phase transitions–among which, those of black holes in anti-de Sitter (AdS) spacetimes have emerged as a focal point of recent research. A core assumption underpinning these studies is the interpretation of the cosmological constant $\Lambda$ as a thermodynamic pressure variable (Kastor:2009wy, ; Dolan:2010ha, ; Kubiznak:2016qmn, ), defined as
@@
where $L$ denotes the AdS radius and $G_{\rm N}$ is Newton’s gravitational constant. Within the framework of this definition, a variety of distinct phase transitions emerge in black hole thermodynamics, such as van der Waals Kubiznak:2012wp , (multiple) re-entrant Altamirano:2013ane ; Frassino:2014pha ; Dehyadegari:2017hvd ; Hendi:2017mfu , those with isolated critical points Dolan:2014vba , those with triple points Altamirano:2013uqa , and superfluid-like phase transitions Hennigar:2016xwd . (See Refs.Bhattacharya:2017hfj ; Spallucci:2013osa ; Majhi:2016txt ; Altamirano:2014tva ; Cai:2013qga ; Hu:2018qsy for more details on related works and extended analyses of these phase transitions.) [^1] Notably, recent studies have revealed that scaling laws in black hole systems can be violated in four dimensions Hu:2024ldp and higher dimensions Ahmed:2022kyv . For the former case, thermodynamic curvature in phase transitions has also been investigated to probe the underlying microstructural properties Abdusattar:2024zzi .
@@
By analyzing the phase transitions of black holes, thermodynamic geometry has emerged as a powerful tool for probing their microscopic structure, with Ruppeiner geometry playing a particularly pivotal role (Ruppeinerb2008, ). The core idea of this approach is to define a line element in the thermodynamic parameter space, which quantifies the “distance” between two adjacent fluctuation states (Ruppeiner95, ). The curvature scalar constructed from this line element directly indicates the nature of interactions among the system’s components: a positive curvature corresponds to repulsive interactions, while a negative curvature signifies attractive interactions. Notably, this interpretation is not unique to black holes, but is derived from the application of Ruppeiner geometry to ordinary thermodynamic systems (Janyszek1990, ; Oshima1999, ), whose cross-system consistency lends reliability to microscopic studies of black holes. Inspired by this, the microscopic origin of black hole thermodynamics has been explored Wei:2012ui ; Wei:2015iwa ; Wei:2019uqg ; Wei:2019yvs ; Guo:2019oad ; Miao:2017fqg ; KordZangeneh:2017lgs ; Wei:2019ctz ; Dehyadegari:2016nkd ; Hu:2020pmr ; Abdusattar:2023xxs . It is important to note that the statistical mechanical study of black holes differs fundamentally from that of ordinary thermodynamic systems: while ordinary systems typically derive macroscopic thermodynamic quantities from the behavior of microscopic particles, black hole research requires “reverse deduction”–inferring information about microscopic structures from known thermodynamic quantities by analyzing macroscopic phase transition behavior.
@@
This paper focuses on the Hayward-AdS black hole, investigating its microstructure through a $P$-$V$ criticality analysis. A well-recognized limitation of classical gravity theories is the existence of spacetime singularities; quantum effects are widely believed to remedy this defect, highlighting the necessity of developing a quantum theory of gravity. However, no mature quantum gravity theory currently exists, making regular black hole models within a semiclassical framework an important avenue for studying singularity correction. Bardeen was the first to propose a singularity free regular black hole model (bardeen1968non, ), which replaces the singularity of traditional black holes with a de Sitter core. Subsequent studies further demonstrated that the charged version of the Bardeen regular black hole can be derived from Einstein gravity coupled to a nonlinear electrodynamic source (AyonBeato:1998ub, ; AyonBeato:2000zs, ), and can even be interpreted as the gravitational field generated by a nonlinear magnetic monopole. Hayward proposed another representative class of regular black hole solutions (Hayward:2005gi, ): similar to the Bardeen solution, it is a degenerate configuration of the gravitational field of a nonlinear source, carrying a magnetic charge and a free integration constant. In theories with infinite towers of higher-order curvature corrections in dimensions $D\geq 5$, regular black holes with anti-de Sitter asymptotics are constructed Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025yqm ; the regularity of these black holes, along with related branes and solutions coupled to various electrodynamics, has been analyzed. Ref. Hennigar:2025yqm also investigates the thermodynamics of such AdS black holes, revealing thermodynamic features associated with regularization parameters. Motivated by these works, this paper focuses on analyzing the phase structure of the Hayward-AdS black hole and exploring the repulsive interactions among its microscopic components.
@@ II Thermodynamics of the Hayward-AdS black holes
Recent advances demonstrate that infinite towers of higher-order curvature corrections to the Einstein-Hilbert action generically resolve the Schwarzschild singularity in spacetimes of dimension $D\geq 5$ Bueno:2024dgm . These corrections underpin the framework of quasi-topological gravities Oliva:2010eb –a class of vacuum gravitational effective field theories featuring second-order field equations in spherical symmetry and a Birkhoff theorem that guarantees the uniqueness of static solutions. This singularity resolution via resummation of infinite curvature corrections aligns with quantum gravity predictions, and while quasi-topological gravities include Lovelock gravity Lovelock:1971yv (which has second-order equations universally), they extend beyond the Lovelock class (limited to $n\leq D/2$):starting with a cubic theory Myers:2010ru , extensions to quartic Dehghani:2011vu , quintic Cisterna:2017umf and all curvature orders Bueno:2019ycr ; Bueno:2022res have been found, all satisfying the Birkhoff theorem such that general vacuum spherically symmetric solutions are static Bueno:2024zsx .
The action for quasi-topological gravity at every curvature order (except for $n=D/2$) and in dimensions $D\geq 5$ is given by Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025yqm
@@
where $g$ is the metric determinant, $R$ is the Ricci scalar, ${\alpha}_{n}$ (length2(n-1)) are coupling constants, and $\mathscr{Z}_{n}$ denotes the nthn-th order curvature term. Here, $\mathscr{Z}_{1}$ corresponds to the Einstein-Hilbert term, $\mathscr{Z}_{2}$ to the Gauss-Bonnet density, and higher-order terms no longer overlap with Lovelock densities. Unlike Lovelock theories, quasi-topological gravities allow infinite curvature corrections–a critical distinction enabling singularity resolution Bueno:2024dgm .
Varying the action in Eq.(2) yields the field equations, from which the line element of the static, spherically symmetric Hayward-AdS black hole is given by Hennigar:2025yqm
@@
Here, ${\rm d}\Omega^{2}_{D-2}$ represents the line element of a (D2)(D-2)-dimensional sphere, $m$ is an integration constant that is related to the Arnowitt-Deser-Misner (ADM) mass or thermodynamic mass $M$ Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025yqm
@@
where $\Omega_{D-2}$ is the volume of a unit sphere $2\pi^{(D-1)/2}/\Gamma\left[(D-1)/2\right]$. The event horizon of this black hole is determined by the condition $f(r_{+})=0$, using which the explicit expression for mass $M$ can be written as Hennigar:2025yqm
$M=\frac{(D-2)\Omega_{D-2}r_{+}^{D-1}}{16\pi G_{\rm N}L^{2}}\Big(1+\frac{L^{2}}{r_{+}^{2}-\alpha}\Big)\,.$ (6)
The Hawking temperature of the Hayward-AdS black hole on an event horizon is obtained as Hennigar:2025yqm
@@
the first law of black hole thermodynamics and the corresponding Smarr formula take the form, respectively Dolan:2014vba ; Hennigar:2025yqm
@@ III.1 $P$-$V$ phase transition and critical behavior
By rearranging the expression in Eq. (7) and combining it with Eq. (1), one can derive the thermodynamic equation of state for the $D$-dimensional spherically symmetric Hayward-AdS black hole, which is given by Hennigar:2025yqm
@@
has a critical point solution $T=T_{\rm c},\ P=P_{\rm c},\ V=V_{\rm c}$. This system is known to exhibit a small-large black hole phase transition Hennigar:2025yqm , where the critical point
@@
The corresponding phase diagram $\widetilde{P}-$$\widetilde{V}$ is presented in Fig. 1.
Figure 1: $\widetilde{P}-$$\widetilde{V}$ phase diagram of Hayward-AdS black hole for different temperatures with $\alpha=0.2,G_{\rm N}=1$.
As depicted in Fig. 1, when the temperature is below the critical value ($T<T_{c}$), oscillatory behavior emerges in the $\widetilde{P}-$$\widetilde{V}$ isotherms. This behavior manifests the coexistence of two distinct phases, namely small and large black holes, and thus signals a first-order phase transition between these two black hole phases analogous to the liquid-gas phase transition in a van der Waals fluid. At critical temperature ($T=T_{c}$), as represented by the red curves, the oscillatory feature in the plots $\widetilde{P}-$$\widetilde{V}$ vanishes, implying that a second-order phase transition occurs precisely at the critical point $(\widetilde{V},\widetilde{P})=(1,1)$. For temperatures above the critical temperature ($T>T_{c}$), the black hole remains in a single thermodynamic phase and does not exhibit any phase transition.
@@ III.2 Critical exponents near the critical point
where $t=\widetilde{T}-1$, $\omega=\widetilde{V}-1$. These two reduced parameters characterize the critical behavior of the system in the vicinity of the critical point. Here, ‘$s$’ and ‘$l$’ are merely subscripts of $V_{s}$ and $V_{l}$, standing for ‘small’ and ‘large’, respectively. Since the black hole entropy is also determined by its horizon radius or volume, $C_{V}$ (constant-volume heat capacity) vanishes–implying the first critical exponent $\widetilde{\alpha}=0$. To calculate the remaining three critical exponents, it is convenient to expand the equation of state given in Eq. (12) around the critical point
@@
$a_{10}$ $=$ $\frac{(D-2){T_{c}}\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{3}{D-1}}}{4G_{\rm N}{P_{c}}\left\{\alpha-\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{2}{D-1}}\right\}^{2}}>0\,$, (21)
$a_{11}$ $=$ $-\frac{\left\{(D-2){T_{c}}\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{3}{D-1}}\left[3\alpha+\left(\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right)^{\frac{2}{D-1}}\right]\right\}}{4\left\{(D-1)G_{\rm N}{P_{c}}\left[\left(\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right)^{\frac{2}{D-1}}-\alpha\right]^{3}\right\}}\neq 0\,$,\,\,
@@
$-D\left(D(D(D(D(41-19D)+114)-333)+189)+27\sqrt{3}\sqrt{(D-2)D}\right)\,$,
$\Theta$ $=$ $\left(2\sqrt{3}\sqrt{(D-2)D^{5}}+3D^{3}-27D-9\sqrt{3}\sqrt{(D-2)D}+27\right)\,$,
all of which are not zero. It is straightforward to show that since $\alpha$ is positive, $a_{11}$ cannot be zero and thus does not affect the critical exponent, which we demonstrate in the following discussions. From Eq.(20), we derive $d\widetilde{P}=(a_{11}t+3a_{03}\omega^{2})d\omega$, with $t$ constant along the isothermal curve. Below the critical temperature ($i.e.,$ $T<T_{c}$), the black hole system exhibits two coexisting phases: the small black hole phase and the large black hole phase. We thus denote the solutions corresponding to these phases as ($\omega_{s}$, $\widetilde{V}_{s}$) (small black hole) and ($\omega_{l}$, $\widetilde{V}_{l}$) (large black hole).
Substituting this into the Maxwell equal-area law Spallucci:2013osa ; Majhi:2016txt ; Bhattacharya:2017hfj ; Lan:2015bia -given by ${P}^{\ast}\left(\widetilde{V}_{s}-\widetilde{V}_{l}\right)=\int^{s}_{l}\widetilde{P}d\widetilde{V}$, we further obtain
@@
where ${P}^{\ast}$ denotes the reduced pressure corresponding to the straight line connecting the two coexisting phases. Additionally, the endpoints of the small and large black hole phases (analogous to the vapor and liquid phases in classical systems) share the same pressure $i.e.,$ ${P}^{\ast}=\widetilde{P}_{l}=\widetilde{P}_{s}$ Spallucci:2013osa ; Majhi:2016txt ; Bhattacharya:2017hfj ; Lan:2015bia . This implies
@@
To conclude, we derive four critical exponents for the Hayward-AdS black hole, given by
@@
These exponents are identical to those of the van der Waals liquid-gas system and thus satisfy the thermodynamic scaling laws predicted by mean field theory Kubiznak:2012wp , as follows